OpenStudy (anonymous):

I am given an interger slope and y-intercept of -49.95-20.00x however I can not find the correct number for the y-intercept can you help me please

OpenStudy (mathmale):

Good morning, Sondra! I'll take your expression (-49.95-20.00x) and write it as the equation of a straight line in slope-intercept form: y = -49.95-20.00x or y = -20.00x - 49.95 compare this to y = m* x + b. What is the value of b here? That's your y-intercept. You could also write the y-intercept as (0,b) (a point).